OpenStudy (anonymous):

Can someone show me how to factor??? Here is the problem: 6z^4-24z^2

OpenStudy (anonymous):

6z^2(z^2-4)

OpenStudy (anonymous):

6z^2(z^2-4) 6z^2(z-2)(z+2)

OpenStudy (anonymous):

take a common factor 6z^2

OpenStudy (saifoo.khan):

6z^4-24z^2 6z^2 is common so, 6z^2(z^2-4) 6z^2(z-2)(z+2)

OpenStudy (anonymous):

then 6z^2(z^2-4) = 6z^2(z-2)(z+2)

OpenStudy (anonymous):

okay and why do you have to put (z-2)(z+2)?

OpenStudy (anonymous):

at the end?

OpenStudy (anonymous):

because you have to factor \[z^2-4\]

OpenStudy (amistre64):

factoring and multiplying are two sides of the same coin

OpenStudy (amistre64):

for example: 3 * 4 = 12 12 is the product of 3 and 4 3 and 4 are factors of 12

OpenStudy (anonymous):

i see and if it was z^2+4 at the you would have to put (z+2)(z+2)? right?

OpenStudy (amistre64):

youre trying to make a pattern; that is not how you factor

OpenStudy (amistre64):

z^2 + 4 doesnt factor across the "real" nukbers

OpenStudy (amistre64):

or numbers even ....

OpenStudy (amistre64):

there are some patterns that emerge, but they are not universal in their applications

OpenStudy (amistre64):

the difference of 2 squares factors into conjugates

OpenStudy (amistre64):

(a^2 - b^2) = (a+b)(a-b)

OpenStudy (amistre64):

the sum of 2 squares does not factor like that

OpenStudy (anonymous):

oh nvm...how would you factor 49x^2-81?

OpenStudy (amistre64):

notice that it is a difference of 2 squares, and follow the pattern for it :)

OpenStudy (amistre64):

(49x^2 - 81) [(7x)^2 - 9^2] = (7x-9)(7x+9)

OpenStudy (anonymous):

okay i think i'm starting to understand, but please please please PLEASE stay in case i need more help.

OpenStudy (amistre64):

i got class starting real soon so i cant make any promises

OpenStudy (anonymous):

okay you can go to class., i got it! =D

OpenStudy (anonymous):

thank you for your help:)